# Riemannian Geometry of Optimal Rebalancing in Dynamic Weight Automated Market Makers

### 05 Mar 2026 v1

@@ 2.1 Dynamic Weight Automated Market Makers
Consider an $N$-token G3M pool with weights $\boldsymbol{\mathbf{w}}=(w_{1},\ldots,w_{N})$ on the simplex ($\sum_{i}w_{i}=1$, $w_{i}>0$) and reserves $\boldsymbol{\mathbf{R}}\in\mathbb{R}^{N}_{+}$. The pool invariant is $\prod_{i}R_{i}^{w_{i}}=k$ [4, 5]. When the pool is in equilibrium, its quoted prices match market prices.
@@ 3 Arbitrage loss is a KL divergence
[Retention Ratio] For an $N$-asset G3M pool with weights changing from $\boldsymbol{\mathbf{w}}^{\mathrm{start}}$ one block to $\boldsymbol{\mathbf{w}}^{\mathrm{end}}$ the next block, with zero fees, perfect arbitrage and, constant prices, the ratio of initial value to final value of the pool is the retention ratio
@@ 5 SLERP midpoint equals (AM+GM)/normalise midpoint
For any $N$-token pool with start weights $\boldsymbol{\mathbf{w}}^{\mathrm{start}}$ and end weights $\boldsymbol{\mathbf{w}}^{\mathrm{end}}$, the SLERP midpoint ($\boldsymbol{\eta}(t)\leftarrow\dfrac{\sin\!\left((1-t)\,\Omega\right)}{\sin\Omega}\,\boldsymbol{\eta}^{\mathrm{start}}+\dfrac{\sin\!\left(t\,\Omega\right)}{\sin\Omega}\,\boldsymbol{\eta}^{\mathrm{end}}$ at $t=k/f=1/2$) is exactly equal to the (AM+GM)/normalise midpoint of [1].
@@ 6 Bounding SLERP’s sub-optimality on the exact KL cost
where $A$ depends on the weight configuration and $\epsilon$ but not on $f$. The relative sub-optimality is $O(\Omega^{2}/f^{2})$.
@@ 8 Discussion
The Lambert W approach of [1] optimises each weight component independently and renormalises, which is only exact for small weight changes. SLERP operates intrinsically on the sphere and automatically preserves the constraint $\sum_{i}w_{i}=1$ (since $\sum_{i}\eta_{i}^{2}=1$ along any great circle in the positive orthant), making it a natural $N$-token generalisation with no renormalisation required.
@@
Computing SLERP requires one $\arccos$ evaluation, $N$ square roots, and per-step $\sin$ calls, comparable in cost to geometric mean interpolation. For on-chain implementations where transcendental functions are expensive, the (AM+GM)/normalise heuristic of [1] remains the most practical choice. The recursive bisection algorithm (Corollary 3) shows that this heuristic computes exact geodesic points at power-of-two step counts using only elementary arithmetic. No trigonometric functions are needed.
@@ Appendix C Deferred proof: Theorem 3 [sub-optimality bound]
The Hessian of $Q$ restricted to the simplex constraints is block-diagonal in the intermediate weight vectors. Each diagonal block receives contributions from two consecutive Fisher–Rao metrics $\mathrm{diag}(1/w_{i}^{(k-1)})$ and $\mathrm{diag}(1/w_{i}^{(k)})$ (the latter from step k+1k{+}1, where $w^{(k)}$ is the base point). On the simplex hyperplane $\{\sum_{i}v_{i}=0\}$, each Fisher–Rao metric has minimum eigenvalue $\geq 1/\max_{i}w_{i}\geq 1$, so each block satisfies $\lambda_{\min}\geq 2$. The per-step KL cost is proportional to $(1{+}u)\log(1{+}u)$, whose second derivative is 1/(1+u)1/(1{+}u), while the quadratic model has second derivative $1$. The Hessian perturbation per block is therefore proportional to 1/(1+u)1=u/(1+u)1/(1{+}u)-1=-u/(1{+}u), which satisfies $|u/(1{+}u)|\leq|u|/(1-|u|)\leq 2|u|$ for $|u|\leq 1/2$. By the Step 1 bound $|u|\leq 2\Omega/(f\epsilon)$, the perturbation magnitude is at most $4\Omega/(f\epsilon)$ per block. For $f\geq 4\Omega/\epsilon$ (already required in Step 2) this is at most $1$, so the Hessian of $C$ satisfies $\lambda_{\min}(\nabla^{2}C)\geq 2-1=1$ per block.

### 09 Mar 2026 v2

@@ 2.1 Dynamic Weight Automated Market Makers
Consider an NN-token G3M pool with weights $\boldsymbol{\mathbf{w}}=(w_{1},\ldots,w_{N})$ on the simplex ($\sum_{i}w_{i}=1$, $w_{i}>0$) and reserves $\boldsymbol{\mathbf{R}}\in\mathbb{R}^{N}_{+}$. The pool invariant is $\prod_{i}R_{i}^{w_{i}}=k$ [4, 5]. When the pool is in equilibrium, its quoted prices match market prices.
@@ 3 Arbitrage loss is a KL divergence
[Retention Ratio] For an NN-asset G3M pool with weights changing from $\boldsymbol{\mathbf{w}}^{\mathrm{start}}$ one block to $\boldsymbol{\mathbf{w}}^{\mathrm{end}}$ the next block, with zero fees, perfect arbitrage and, constant prices, the ratio of initial value to final value of the pool is the retention ratio
@@ 5 SLERP midpoint equals (AM+GM)/normalise midpoint
For any NN-token pool with start weights $\boldsymbol{\mathbf{w}}^{\mathrm{start}}$ and end weights $\boldsymbol{\mathbf{w}}^{\mathrm{end}}$, the SLERP midpoint ($\boldsymbol{\eta}(t)\leftarrow\dfrac{\sin\!\left((1-t)\,\Omega\right)}{\sin\Omega}\,\boldsymbol{\eta}^{\mathrm{start}}+\dfrac{\sin\!\left(t\,\Omega\right)}{\sin\Omega}\,\boldsymbol{\eta}^{\mathrm{end}}$ at $t=k/f=1/2$) is exactly equal to the (AM+GM)/normalise midpoint of [1].
@@ 6 Bounding SLERP’s sub-optimality on the exact KL cost
where AA depends on the weight configuration and $\epsilon$ but not on $f$. The relative sub-optimality is $O(\Omega^{2}/f^{2})$.
@@ 8 Discussion
The Lambert W approach of [1] optimises each weight component independently and renormalises, which is only exact for small weight changes. SLERP operates intrinsically on the sphere and automatically preserves the constraint $\sum_{i}w_{i}=1$ (since $\sum_{i}\eta_{i}^{2}=1$ along any great circle in the positive orthant), making it a natural NN-token generalisation with no renormalisation required.
@@
Computing SLERP requires one $\arccos$ evaluation, NN square roots, and per-step $\sin$ calls, comparable in cost to geometric mean interpolation. For on-chain implementations where transcendental functions are expensive, the (AM+GM)/normalise heuristic of [1] remains the most practical choice. The recursive bisection algorithm (Corollary 3) shows that this heuristic computes exact geodesic points at power-of-two step counts using only elementary arithmetic. No trigonometric functions are needed.
@@ Appendix C Deferred proof: Theorem 3 [sub-optimality bound]
The Hessian of $Q$ restricted to the simplex constraints is block-diagonal in the intermediate weight vectors. Each diagonal block receives contributions from two consecutive Fisher–Rao metrics $\mathrm{diag}(1/w_{i}^{(k-1)})$ and $\mathrm{diag}(1/w_{i}^{(k)})$ (the latter from step k+1k{+}1, where $w^{(k)}$ is the base point). On the simplex hyperplane $\{\sum_{i}v_{i}=0\}$, each Fisher–Rao metric has minimum eigenvalue $\geq 1/\max_{i}w_{i}\geq 1$, so each block satisfies $\lambda_{\min}\geq 2$. The per-step KL cost is proportional to $(1{+}u)\log(1{+}u)$, whose second derivative is 1/(1+u)1/(1{+}u), while the quadratic model has second derivative 11. The Hessian perturbation per block is therefore proportional to 1/(1+u)1=u/(1+u)1/(1{+}u)-1=-u/(1{+}u), which satisfies $|u/(1{+}u)|\leq|u|/(1-|u|)\leq 2|u|$ for $|u|\leq 1/2$. By the Step 1 bound $|u|\leq 2\Omega/(f\epsilon)$, the perturbation magnitude is at most $4\Omega/(f\epsilon)$ per block. For $f\geq 4\Omega/\epsilon$ (already required in Step 2) this is at most 11, so the Hessian of $C$ satisfies $\lambda_{\min}(\nabla^{2}C)\geq 2-1=1$ per block.